# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v2

@@ A.3 Proof of Theorem 4
$\mathcal{E}(t)\leq\alpha(t)\exp{\int_{t}^{\tau_{k}}\gamma(s)}\mathrm{d}s.$ (44)
Therefore, the reconstruction error of HOG-Diff is bounded by
$\begin{split}\mathcal{E}(0)&\leq\alpha(0)\exp{\int_{0}^{\tau_{1}}\gamma(s)}\mathrm{d}s\\ &=\left(C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathrm{d}s+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathrm{d}s\right)\exp{\int_{0}^{\tau_{1}}\gamma(s)}\mathrm{d}s.\end{split}$ (45)
A comparable calculation for a classical graph generation model (with diffusion interval $[0,T]$) yields a bound
$\mathcal{E}^{\prime}(0)\leq\left(C^{2}\ell(\bm{\varphi})\int_{0}^{T}g_{s}^{4}\mathrm{d}s+C\mathcal{E}^{\prime}(T)\int_{0}^{T}h_{s}^{2}\mathrm{d}s\right)\exp\int_{0}^{T}\gamma^{\prime}(s)\mathrm{d}s,$ (46)
where $h_{s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:T}}-1}\right)$.
Let $h(s,\tau)=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau}}-1}\right)$, $a(\tau)=\int_{0}^{\tau}h(s,\tau)^{2}\mathop{}\!\mathrm{d}{s}$, and $b(\tau)=\int_{0}^{\tau}\|h(s,\tau)\|_{\operatorname{lip}}^{2}\mathop{}\!\mathrm{d}{s}$ . Since $\tau_{1}\leq T$, it follows that $\mathcal{E}(\tau_{1})\leq\mathcal{E}^{\prime}(T)$. Additionally, by Theorem 3, $\ell(\cdot)$ converges exponentially in the score-matching process. Therefore, to prove $\mathcal{E}(0)\leq\mathcal{E}^{\prime}(0)$, it suffices to show that both $a(\tau)$ and $b(\tau)$ are increasing functions.
Applying the Leibniz Integral Rule, we obtain:
$a^{\prime}(\tau)=h(\tau,\tau)^{2}+\int_{0}^{\tau}\frac{\partial}{\partial\tau}h(s,\tau)^{2}\mathop{}\!\mathrm{d}{s}\qquad\mathrm{and}\qquad b^{\prime}(\tau)=\|h(\tau,\tau)\|^{2}_{\operatorname{lip}}+\int_{0}^{\tau}\frac{\partial}{\partial\tau}\|h(s,\tau)\|^{2}_{\operatorname{lip}}\mathop{}\!\mathrm{d}{s}.$ (47)
Since $h(\tau,\tau)\rightarrow 0$, we can derive that $a^{\prime}(\tau)>0$ and $b^{\prime}(\tau)>0$. This implies $\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}\leq\int_{0}^{T}h_{s}^{2}\mathop{}\!\mathrm{d}{s}$ and $\int_{0}^{\tau_{1}}\gamma(s)\mathop{}\!\mathrm{d}{s}\leq\int_{0}^{T}\gamma^{\prime}(s)\mathop{}\!\mathrm{d}{s}$. Combining these inequalities, we can finally conclude $\mathcal{E}(0)\leq\mathcal{E}^{\prime}(0)$. Therefore, HOG-Diff provides a sharper reconstruction error bound than classical graph generation frameworks. ∎
@@ B.1 Higher-order Networks
Graphs are elegant and useful abstractions for modeling irregular relationships in empirical systems, transforming unstructured data into analyzable representations. However, their inherent limitation to pairwise interactions restricts their representation of group dynamics (Battiston et al., 2020). For example, cyclic structures like benzene rings and functional groups play a holistic role in molecular networks; densely interconnected structures, like simplices, often have a collective influence on social networks; and functional brain networks exhibit higher-order dependencies. To address this, various topological models have been employed to describe data in terms of its higher-order relations, including simplicial complexes, cell complexes, and combinatorial complexes (Hajij et al., 2023). As such, the study of higher-order networks has gained increasing attention for their capacity to capture higher-order interactions, with broad applications across domains such as social network analysis (Zeng et al., 2024), graph signal processing (Roddenberry et al., 2022; Sardellitti et al., 2021), and topological deep learning (Bodnar et al., 2021; Huang et al., 2024; Papamarkou et al., 2024).
Given the broad applicability and theoretical richness of higher-order networks, the following delves deeper into two key frameworks for modelling such interactions: simplicial complexes and cell complexes.
@@ B.2 Graph Generative Models
The study of graph generation seeks to synthesize graphs that align with the observed distribution. Graph generation has been extensively studied, which dates back to the early works of the random network models, such as the Erdős–Rényi (ER) model (Erdős et al., 1960) and the Barabási-Albert (BA) model (Barabási & Albert, 1999). While these models offer foundational insights, they are often too simplistic to capture the complexity of graph distributions we encounter in practice.
Recent graph generative models have made great progress in graph distribution learning by exploiting the capacity of deep neural networks. GraphRNN (You et al., 2018) and GraphVAE (Simonovsky & Komodakis, 2018) adopt sequential strategies to generate nodes and edges. MolGAN (De Cao & Kipf, 2018) integrates generative adversarial networks (GANs) with reinforcement learning objectives to synthesize molecules with desired chemical properties. Shi et al. (2020) generates molecular graphs using a flow-based approach, while GraphDF (Luo et al., 2021) adopts an autoregressive flow-based model with discrete latent variables. Additionally, GraphEBM (Liu et al., 2021) employs an energy-based model for molecular graph generation. However, the end-to-end structure of these methods often makes them more challenging to train compared to diffusion-based generative models.
@@
Diffusion bridge processes, i.e., processes conditioned to the endpoints, have been widely adopted in image-related domains, including image generation (De Bortoli et al., 2021), image translation (Zhou et al., 2024), and image restoration (Luo et al., 2023b; Yue et al., 2024). Recently, several studies have improved the graph generative framework of diffusion models by leveraging the diffusion bridge processes. Wu et al. (2022) inject physical information into the process by incorporating informative prior to the drift. GLAD (Boget et al., 2024) employs the Brownian bridge on a discrete latent space with endpoints conditioned on data samples. GruM (Jo et al., 2024) utilizes the OU bridge to condition the diffusion endpoint as the weighted mean of all possible final graphs. However, existing methods often overlook or inadvertently disrupt the higher-order topological structures in the graph generation process.
@@ C.2 Spectral Diffusion
Classical graph diffusion approaches typically inject isotropic Gaussian noise directly into adjacency matrices $\bm{A}$, leading to various fundamental challenges. Firstly, the inherent non-uniqueness of graph representations implies that a graph with $n$ vertices can be equivalently modelled by up to $n!$ distinct adjacency matrices. This ambiguity requires a generative model that assigns probabilities uniformly across all equivalent adjacencies to accurately capture the graph’s inherent symmetry. Additionally, unlike densely distributed image data, graphs typically follow a Pareto distribution and exhibit sparsity (Ghavasieh & De Domenico, 2024), so that adjacency score functions lie on a low-dimensional manifold. Consequently, noise injected into out-of-support regions of the full adjacency space severely degrades the signal-to-noise ratio, impairing the training of the score-matching process. Even for densely connected graphs, isotropic noise distorts global message-passing patterns by encouraging message-passing on sparsely connected regions. Moreover, the adjacency matrix scales quadratically with the number of nodes, making the direct generation of adjacency matrices computationally prohibitive for large-scale graphs.
@@
Here, the superscript of $\bm{X}^{(k)}_{t}$ and $\bm{\Lambda}^{(k)}_{t}$ are dropped for simplicity, and $\mathbf{f}_{k,t}$ is determined according to Eq. 6.
@@ C.4 Training and Sampling Procedure
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ controls the relative importance of vertices and spectrum.
@@
Appendix D Complexity Analysis
@@
For simplicial complexes, the number of $p$-simplices in a graph with $n$ nodes and $m$ edges is upper-bounded by $\mathcal{O}(n^{p-1})$, and they can be enumerated in $\mathcal{O}(a\left(\mathcal{G}\right)^{p-3}m)$ time (Chiba & Nishizeki, 1985), where $a\left(\mathcal{G}\right)$ is the arboricity of the graph $\mathcal{G}$, a measure of graph sparsity. Since arboricity is demonstrated to be at most $\mathcal{O}(m^{1/2})$ and $m\leq n^{2}$, all $p$-simplices can thus be listed in $\mathcal{O}\left(n^{p-3}m\right)$. Besides, the complexity of finding $2$-simplex is estimated to be $\mathcal{O}(\left\langle k\right\rangle m)$ with the Bron–Kerbosch algorithm (Bron & Kerbosch, 1973), where $\left\langle k\right\rangle$ denotes the average node degree, typically a small value for empirical networks.
@@
In short, the only scaling cost is a one-off preprocessing pass; both training and sampling remain GPU-bound and on par with, or even faster than, existing diffusion models.

### v3

@@ A.3 Proof of Theorem 4
$\mathcal{E}(t)\leq\alpha_{k}(t)\exp{\int_{t}^{\tau_{k}}\gamma_{k}(s)}\mathrm{d}s.$ (50)
Evaluating the bound at the beginning of the $k$-th reverse stage, that is $t=\tau_{k-1}$, yields
$\mathcal{E}(\tau_{k-1})$ $\leq\alpha_{k}(\tau_{k-1})\exp\left(\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s\right)$ (51)
$=\left[C^{2}\ell^{(k)}(\bm{\theta})\int_{\tau_{k-1}}^{\tau_{k}}g_{k,s}^{4}\mathrm{d}s+C\mathcal{E}(\tau_{k})\int_{\tau_{k-1}}^{\tau_{k}}h_{k,s}^{2}\mathrm{d}s\right]\exp\left(\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s\right).$
For later convenience, we introduce the nonnegative stage-wise coefficients
$A_{k}=C^{2}\ell^{(k)}(\bm{\theta})\int_{\tau_{k-1}}^{\tau_{k}}g_{k,s}^{4}\mathrm{d}s,\quad D_{k}=C\int_{\tau_{k-1}}^{\tau_{k}}h_{k,s}^{2}\mathrm{d}s,\quad T_{k}=\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s.$ (52)
Then the error bound can be written compactly as
$\mathcal{E}(\tau_{k-1})\leq\bigl(A_{k}+D_{k}\mathcal{E}(\tau_{k})\bigr)\,e^{T_{k}},\qquad k=1,\cdots,K.$ (53)
The classical diffusion model corresponds to a single reverse stage on $[0,T]$, and its error bound can be formulated in the same functional form as
$\mathcal{E}_{\mathrm{cls}}(0)\leq A_{\mathrm{cls}}\,e^{T_{\mathrm{cls}}},$ (54)
where $A_{\mathrm{cls}}:=C^{2}\ell_{\mathrm{cls}}(\bm{\varphi})\int_{0}^{T}g_{s}^{4}\,\mathrm{d}s$ and $T_{\mathrm{cls}}:=\int_{0}^{T}\gamma_{\mathrm{cls}}(s)\mathrm{d}s+C\int_{0}^{T}h_{s}^{2}\mathrm{d}s$.
Step 2: Multi-stage accumulated error bound.
The reverse process starts from the same prior for both oracle and learned SDEs, hence $\mathcal{E}(\tau_{K})=0$. Unrolling the recursion Eq. 53 backward from $k=K$ yields
$\mathcal{E}(\tau_{K-1})\leq A_{K}e^{T_{K}},$ (55)
and
$\mathcal{E}(\tau_{K-2})\leq\left(A_{K-1}+D_{K-1}\mathcal{E}(\tau_{K-1})\right)e^{T_{K-1}}\leq\left(A_{K-1}+D_{K-1}A_{K}e^{T_{K}}\right)e^{T_{K-1}}.$ (56)
Continuing inductively, one obtains for the initial time $t=0=\tau_{0}$ the general bound
$\mathcal{E}_{\mathrm{hog}}(0)\leq\sum_{k=1}^{K}A_{k}\exp\left(\sum_{m=1}^{k}T_{m}\right)\prod_{j=1}^{k-1}D_{j},$ (57)
where the empty product is defined as $\prod_{j=1}^{0}D_{j}:=1$.
Since each $D_{j}$ is nonnegative, the elementary inequality $D_{j}\leq e^{D_{j}}$ for all $x\geq 0$ implies
$\mathcal{E}_{\mathrm{hog}}(0)\leq\sum_{k=1}^{K}A_{k}\exp\left(\sum_{m=1}^{k}T_{m}+\sum_{j=1}^{k-1}D_{j}\right)\leq\left(\sum_{k=1}^{K}A_{k}\right)\exp\left(\sum_{m=1}^{K}T_{m}+\sum_{j=1}^{K-1}D_{j}\right).$ (58)
The latter inequality holds as all quantities in the exponent are nonnegative, so enlarging the summation ranges can only increase the exponent.
Therefore, we obtain the compact multi-stage error bound
$\mathcal{E}_{\mathrm{hog}}(0)\leq A_{\mathrm{hog}}\,e^{T_{\mathrm{hog}}},$ (59)
where $A_{\mathrm{hog}}:=\sum_{k=1}^{K}A_{k}$, and $T_{\mathrm{hog}}:=\sum_{m=1}^{K}T_{m}+\sum_{j=1}^{K-1}D_{j}$. This bound admits the same functional form as in Eq. 54 but accumulates the contributions of all $K$ reverse stages from $T=\tau_{K}$ to $0=\tau_{0}$.
Step 3: Comparing the two error bounds.
We now compare $(A_{\mathrm{cls}},T_{\mathrm{cls}})$ in Eq. 54 with $(A_{\mathrm{hog}},T_{\mathrm{hog}})$ in Eq. 59.
First, by construction each subinterval $\tau_{k-1},\tau_{k}]$ lies inside $[0,T]$ and the OU bridge horizon of the $k$-th HOG-Diff stage satisfies $\tau_{k}\leq T$. The corresponding drift and diffusion coefficients are therefore pointwise dominated by those of the classical bridge:
$g_{k,s}^{4}\leq g_{s}^{4},\qquad h_{k,s}^{2}\leq h_{s}^{2},\qquad\gamma_{k}(s)\leq\gamma_{\mathrm{cls}}(s),\qquad s\in[\tau_{k-1},\tau_{k}].$ (60)
Since all these quantities are nonnegative, their integrals over a subinterval are bounded by the integrals over the full horizon:
$\sum_{k=1}^{K}\int_{\tau_{k-1}}^{\tau_{k}}g_{k,s}^{4}\mathrm{d}s\leq\int_{0}^{T}g_{s}^{4}\mathrm{d}s,\quad\sum_{k=1}^{K}\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s\leq\int_{0}^{T}\gamma_{\mathrm{cls}}(s)\mathrm{d}s,$ (61)
and
$\sum_{k=1}^{K}\int_{\tau_{k-1}}^{\tau_{k}}h_{k,s}^{2}\mathrm{d}s\leq\int_{0}^{T}h_{s}^{2}\,\mathrm{d}s.$ (62)
For the coefficient $A_{\mathrm{hog}}$, using the definition in Eq. 52 we have
$A_{\mathrm{hog}}=C^{2}\sum_{k=1}^{K}\ell^{(k)}(\bm{\theta})\int_{\tau_{k-1}}^{\tau_{k}}g_{k,s}^{4}\,\mathrm{d}s\leq C^{2}\left(\sum_{k=1}^{K}\ell^{(k)}(\bm{\theta})\right)\int_{0}^{T}g_{s}^{4}\,\mathrm{d}s.$ (63)
At the level of optimal score-matching losses, Sec. A.3 shows that the coarse-to-fine curriculum of HOG-Diff cannot be intrinsically worse than the single-stage formulation: $\sum_{k=1}^{K}\ell^{(k)}_{\star}\leq\ell_{\mathrm{cls},\star}$. In the regime where the optimization error is negligible and the loss is dominated by approximation error, it is natural to focus on solutions for which the curriculum does not increase the total score-matching loss, namely
$\sum_{k=1}^{K}\ell^{(k)}(\bm{\theta})\leq\ell_{\mathrm{cls}}(\bm{\varphi}).$ (64)
This condition is clearly satisfied in our experiments. For $K=2$, the combined spectrum loss of the Coarse and Fine stages, denoted as Coarse+Fine in Fig. 4, is lower than that of the classical model (Classical (Noise)), providing empirical support for the above comparison.
Hence, we can bound $A_{\mathrm{hog}}$ as
$A_{\mathrm{hog}}\leq C^{2}\ell_{\mathrm{cls}}(\bm{\varphi})\int_{0}^{T}g_{s}^{4}\mathrm{d}s=A_{\mathrm{cls}}.$ (65)
For the exponent $T_{\mathrm{hog}}$ in Eq. 59, we use the same domination to obtain
$T_{\mathrm{hog}}$ $=\sum_{k=1}^{K}T_{k}+\sum_{k=1}^{K-1}D_{k}=\sum_{k=1}^{K}\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s+C\sum_{k=1}^{K-1}\int_{\tau_{k-1}}^{\tau_{k}}h_{k,s}^{2}\mathrm{d}s$ (66)
$\leq\int_{0}^{T}\gamma_{\mathrm{cls}}(s)\mathrm{d}s+C\int_{0}^{T}h_{s}^{2}\mathrm{d}s$
$=T_{\mathrm{cls}}.$
Together with the bounds in Eq. 54 and Eq. 59, we obtain
$A_{\mathrm{hog}}e^{T_{\mathrm{hog}}}\leq A_{\mathrm{cls}}e^{T_{\mathrm{cls}}}.$ (67)
Combining these inequalities, we can finally conclude $\mathcal{E}_{\mathrm{hog}}(0)\leq\mathcal{E}_{\mathrm{cls}}(0)$, which shows that the reconstruction error bound of HOG-Diff over the full trajectory $T\to 0$ is no worse, and can be strictly sharper, than the corresponding bound for the classical diffusion model.
@@ B.1 Higher-order Networks
Graphs provide elegant and practical abstractions for modeling irregular relationships in empirical systems, transforming unstructured data into analyzable representations. However, their inherent restriction to pairwise interactions limits their ability to represent group dynamics (Battiston et al., 2020). For example, cyclic structures like benzene rings and functional groups play a holistic role in molecular networks; densely interconnected structures, like simplices, often have a collective influence on social networks; and functional brain networks exhibit higher-order dependencies. To address this, various topological formalisms have been employed to describe data in terms of its higher-order relations, including simplicial complexes, cell complexes, and combinatorial complexes (Hajij et al., 2023). As such, the study of higher-order networks has gained increasing attention for their capacity to capture higher-order interactions, with broad applications across domains such as influence maximization (Zeng et al., 2024), financial system analysis (Mai et al., 2024; Xiao et al., 2024), graph signal processing (Sardellitti et al., 2021; Roddenberry et al., 2022), and graph representation learning (Bodnar et al., 2021; Huang et al., 2024; Papamarkou et al., 2024).
Given their broad applicability and theoretical richness, we next focus on two core frameworks for modelling higher-order interactions: simplicial complexes and cell complexes.
@@ B.2 Graph Generative Models
The study of graph generation seeks to synthesize graphs that align with the observed distribution. Graph generation has been extensively studied, which dates back to the early works of the random network models, such as the Erdős–Rényi (ER) model (Erdős et al., 1960) and the Barabási-Albert (BA) model (Barabási and Albert, 1999). While these models offer foundational insights, they are often too simplistic to capture the complexity of graph distributions we encounter in practice.
Recent graph generative models have made great progress in graph distribution learning by exploiting the capacity of deep neural networks. GraphRNN (You et al., 2018) and GraphVAE (Simonovsky and Komodakis, 2018) adopt sequential strategies to generate nodes and edges. MolGAN (De Cao and Kipf, 2018) integrates generative adversarial networks (GANs) with reinforcement learning objectives to synthesize molecules with desired chemical properties. Shi et al. (2020) generates molecular graphs using a flow-based approach, while GraphDF (Luo et al., 2021) adopts an autoregressive flow-based model with discrete latent variables. Additionally, GraphEBM (Liu et al., 2021) employs an energy-based model for molecular graph generation. However, the end-to-end structure of these methods often makes them more challenging to train compared to diffusion-based generative models.
@@
Diffusion bridge processes, i.e., processes conditioned to the endpoints, have been widely adopted in image-related domains, including image generation (De Bortoli et al., 2021), image translation (Zhou et al., 2024), and image restoration (Luo et al., 2023b; Yue et al., 2024). Recently, several studies have improved the graph generative framework of diffusion models by leveraging the diffusion bridge processes. Wu et al. (2022) inject physical information into the process by incorporating informative prior to the drift. GruM (Jo et al., 2024) utilizes the OU bridge to condition the diffusion endpoint as the weighted mean of all possible final graphs. GLAD (Boget et al., 2025) employs the Brownian bridge on a discrete latent space with endpoints conditioned on data samples. However, existing methods often overlook or inadvertently disrupt the higher-order topological structures in the graph generation process.
@@ C.2 Spectral Diffusion
Classical graph diffusion approaches typically inject isotropic Gaussian noise directly into adjacency matrices $\bm{A}$, leading to various fundamental challenges. Firstly, the inherent non-uniqueness of graph representations implies that a graph with $n$ vertices can be equivalently modelled by up to $n!$ distinct adjacency matrices. This ambiguity requires a generative model that assigns probabilities uniformly across all equivalent adjacencies to accurately capture the graph’s inherent symmetry. Additionally, unlike densely distributed image data, graphs typically follow a Pareto distribution and exhibit sparsity (Ghavasieh and De Domenico, 2024), so that adjacency score functions lie on a low-dimensional manifold. Consequently, noise injected into out-of-support regions of the full adjacency space severely degrades the signal-to-noise ratio, impairing the training of the score-matching process. Even for densely connected graphs, isotropic noise distorts global message-passing patterns by encouraging message-passing on sparsely connected regions. Moreover, the adjacency matrix scales quadratically with the number of nodes, making the direct generation of adjacency matrices computationally prohibitive for large-scale graphs.
@@
Here, the superscript of $\bm{X}^{(k)}_{t}$ and $\bm{\Lambda}^{(k)}_{t}$ are dropped for simplicity, and $\mathbf{f}_{k,t}$ is determined according to Eq. 7. Note that, for molecular datasets, the intermediate endpoint is obtained by cell complex filtering: it preserves coarse higher-order connectivity, but the resulting node features can carry only an approximate atom-type signal. To stabilize training, we therefore add a small masked Gaussian perturbation to its node features during training
@@ C.4 Training and Sampling Procedure
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ control the relative importance of vertices and spectrum.
@@
Appendix D Efficiency and Complexity Analysis
D.1 One-time Preprocessing Complexity
@@
For simplicial complexes, the number of $p$-simplices in a graph with $n$ nodes and $m$ edges is upper-bounded by $\mathcal{O}(n^{p-1})$, and they can be enumerated in $\mathcal{O}(a\left(\mathcal{G}\right)^{p-3}m)$ time (Chiba and Nishizeki, 1985), where $a\left(\mathcal{G}\right)$ is the arboricity of the graph $\mathcal{G}$, a measure of graph sparsity. Since arboricity is demonstrated to be at most $\mathcal{O}(m^{1/2})$ and $m\leq n^{2}$, all $p$-simplices can thus be listed in $\mathcal{O}\left(n^{p-3}m\right)$. Besides, the complexity of finding $2$-simplex is estimated to be $\mathcal{O}(\left\langle k\right\rangle m)$ with the Bron–Kerbosch algorithm (Bron and Kerbosch, 1973), where $\left\langle k\right\rangle$ denotes the average node degree, typically a small value for empirical networks.
D.2 Empirical Efficiency: Preprocessing, Training, and Sampling
To complement the theoretical analysis, we empirically evaluate the scalability of our approach on standard benchmarks of varying scales. As shown in Fig. 8, we record the wall-clock time for the filtering procedure on datasets ranging from small community graphs to large-scale molecular datasets such as ZINC250k and MOSES. The results indicate that the filtering time scales reasonably with data size. Given that this is a one-time preprocessing cost, it confirms that HOG-Diff is computationally feasible for large-scale graph generation tasks.
@@ D.2 Empirical Efficiency: Preprocessing, Training, and Sampling
Table 5: Training time comparison between GDSS and HOG-Diff.
Method Community-small Ego-small Enzymes QM9 ZINC250k
GDSS 13 min 19 min 1.9 h 1.2 h 14.0 h
HOG-Diff 20 min 12 min 0.5 h 3.1 h 6.5 h
@@
Figure 8: Empirical scalability of the filtering procedure. We report the wall-clock time required for the filtering (preprocessing) operation across datasets of varying sizes. The results show that the one-time preprocessing cost remains tractable even for large-scale datasets.
@@
Scalability on Large Benchmarks.
As shown in Fig. 8, the one-time filtering cost remains tractable across large-scale benchmarks, including datasets with substantially more training samples (MOSES) and larger graphs (SBM and GuacaMol). Importantly, this preprocessing is incurred only once; after that, the training loop is identical to standard diffusion models and the sampling overhead remains negligible. Detailed generation results on MOSES, GuacaMol, and SBM are reported in Sec. F.2.
In short, HOG-Diff scales gracefully: the extra cost is confined to a one-time preprocessing step, while the end-to-end runtime is dominated by standard GPU diffusion training and sampling.